OpenStudy (anonymous):

Simplify 13 over the quantity of negative 3 plus 2i.

OpenStudy (anonymous):

\[\frac{ 13 }{ 3+2i }\]

OpenStudy (anonymous):

the 3 is suppose to be negative but I got -3+2i @ganeshie8 am I right? \[\frac{ 13\left( -3+2i \right) }{ -3+2i \left( -3-2i \right)} = \frac{ -39+26i }{ 13 }= -3+2i\]

OpenStudy (solomonzelman):

You mean \(\displaystyle \frac{ 13\left( -3+2i \right) }{\color{red}{(} -3+2i \color{red}{)}\left( -3-2i \right)} = \frac{ -39+26i }{ 13 }= -3+2i\)

OpenStudy (solomonzelman):

yes, what you did is correct

OpenStudy (solomonzelman):

just don't leave those parenthesis, so that we can correctly see what you multiply by... for other than this, that was very well done! Good Job!

OpenStudy (anonymous):

thank you